OpenStudy (anonymous):

θ is any real number. is it possible for sinθ=.25 and cosθ=.5? why or why not?

OpenStudy (anonymous):

so is this the same for this problem? -1 ≤ sin θ ≤ 1

OpenStudy (agent0smith):

^yes. I'm assuming that they're not implying the θ in sinθ is the same θ as in cosθ.

OpenStudy (anonymous):

oh ok thanks

OpenStudy (agent0smith):

-1 ≤ cos θ ≤ 1 this too.

OpenStudy (anonymous):

hm i'm not sure what they mean

OpenStudy (agent0smith):

I think the "θ is any real number" just means θ can be anything necessary to satisfy the sinθ=.25 and cosθ=.5, doesn't have to be the same value (i'm guessing).

OpenStudy (anonymous):

you're probably right. so in that case how can i find the answer?

OpenStudy (agent0smith):

If they do mean that it has to be the same value, you'd have to take the inverse sine for sinθ=.25 and inverse cosine for cosθ=.5

OpenStudy (anonymous):

what would that be?

OpenStudy (anonymous):

@agent0smith is there a formula i should use to find it?

OpenStudy (agent0smith):

Have you used inverse sin (written as sin^-1) and inverse cosine (cos^-1) yet? Maybe arcsin and arccos? If neither of those seem remotely familiar, then I wouldn't worry about them for now.

OpenStudy (anonymous):

yes i think i have i just dont really remember it too clearly

OpenStudy (anonymous):

im just not sure what the answer should be here

OpenStudy (agent0smith):

sinθ=.25 inverse sin of both sides gives θ = arcsin0.25 (do this on your calculator) cosθ = 0.5 inverse cosine of both sides gives θ = arccos 0.5

OpenStudy (whpalmer4):

I would argue that the language "is it possible for sinθ=.25 AND cosθ=.5? Why or why not?" means they want you to determine whether or not there is a value of theta for which both are true.

OpenStudy (agent0smith):

^ yeah that's what i started to think too. The "and" hints at that.

OpenStudy (anonymous):

i got θ=.2527

OpenStudy (whpalmer4):

And a quick glance at the neighborhood unit circle says that for that to be true, .25^2 + .5^2 must = 1^2.

OpenStudy (anonymous):

oh ok i see what you mean

OpenStudy (whpalmer4):

It's not a very interesting question otherwise...

OpenStudy (anonymous):

so i should say that there is or is not a value of theta which is true for both?

OpenStudy (whpalmer4):

Is the equation that I said must be true, true?

OpenStudy (agent0smith):

^correct. "i got θ=.2527" - this would be in radians, not degrees.

OpenStudy (whpalmer4):

remember, this is just that basic trig identity sin^x + cos^x = 1

OpenStudy (agent0smith):

@whpalmer4 Idk if they'd have done trig identities yet (but maybe). The last question she posted was basic, just "is it possible for sinθ=5/4".

OpenStudy (anonymous):

ahh im so confused :/ sorry :(

OpenStudy (whpalmer4):

@agent0smith well, that's not inconsistent whether or not they've had the identities so long as they've had the unit circle.

OpenStudy (whpalmer4):

Erin, have you seen the unit circle showing you the relationship between sin and cos?

OpenStudy (anonymous):

yes

OpenStudy (agent0smith):

Ah, then it probably is asking you if there's a θ that satisfies both cos = 0.25 and sin = 0.5.

OpenStudy (agent0smith):

And you could use what whpalmer4 posted above... sin^2 θ + cos^2 θ =1

OpenStudy (whpalmer4):

Okay, in that case, I think the question is as I suggest: they want to know if there is some angle for which both of those statements are true at the same time. And using the Pythagorean theorem in the unit circle (hypotenuse = 1, cos x = a, and sin x = b), you can see that for there to be such an angle, 0.25^2 + 0.5^2 = 1^2. If that is not true, then there is no such angle.

OpenStudy (whpalmer4):

What is (1/2)^2 + (1/4)^2?

OpenStudy (anonymous):

.3125

OpenStudy (anonymous):

so its not true

OpenStudy (anonymous):

ohhh i understand now. thank you both so much, i really appreciate your help

OpenStudy (whpalmer4):

Good!

OpenStudy (anonymous):

can you please help me with another onee? :)

OpenStudy (whpalmer4):

If you are working in radians, for small values of \(\theta\), \(\sin \theta \approx \theta\). You saw that when you found 0.2527 for \(\theta\) but perhaps you didn't realize it :-)

OpenStudy (anonymous):

oh i understand now, thanks so much